OpenStudy (boxman61):

Factor Completely 12x^3-3x

OpenStudy (saifoo.khan):

3x(2x-1)(2x+1)

OpenStudy (boxman61):

that was scary fast..thanks

OpenStudy (saifoo.khan):

hehe. Welcome. ;)

OpenStudy (anonymous):

3x(2x-1)(2x+1)

OpenStudy (anonymous):

3x(2x-1)(2x+1)

OpenStudy (saifoo.khan):

Everyone's copying my answer. lol

OpenStudy (anonymous):

I know but I love factoring so I wanted to try it out. I honestly didnt even look at your answer I did it myself. :)

OpenStudy (saifoo.khan):

Lol, i was just kidding. nice work though. ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

12x³ - 3x Looking at the terms, we see that they both have a 3 in common. Taking it out gives us: 3(4x³ - x) Looking again, we can still take an x out! 3x(4x² - 1) Since 4x² - 1 is in the form of (x² - y²), it can be factored easily (this is called a difference of squares). The rule is x² - y² = (x+y)(x-y). That means that: 4x² - 1 = (2x + 1)(2x - 1) Replacing it back into the expression, 12x³ - 3x = 3x(2x + 1)(2x - 1)